OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x))

OpenStudy (anonymous):

just plug in where the x is at, the whole expression in right side.

OpenStudy (unklerhaukus):

\[f(x)=3(x)+2\] \[g(x)=(x)^2\] \[(f\circ g)(x)=f(g(x))=f((x)^2)=\dots\]

OpenStudy (anonymous):